A metal bar with an initial temperature, 𝑇0, in the interval of 30Β°C ≀ 𝑇0 ≀ 35Β°C is dropped into a container of boiling water (100Β°C). The temperature of the metal bar, 𝑇 at any time, 𝑑 satisfies the following Newton’s Law of Cooling model 𝑑𝑇 𝑑𝑑 = βˆ’π‘˜(𝑇 βˆ’ π‘‡π‘š) where π‘‡π‘š is the ambient temperature and π‘˜ is the constant. After 5 seconds, the temperature of the bar, 𝑇1 is in the interval of 40Β°C ≀ 𝑇1 ≀ 50Β°C. a. Find the equation that models the temperature of the metal bar, 𝑇 at any time, 𝑑 (choose a value of 𝑇0 and 𝑇1 from the given intervals, respectively). By using an appropriate analytical method, solve the derived model and explain the reason for the selection of the method. b. Compute the temperature of the metal bar after 100 seconds by using the derived model in Part 1(a) with THREE (3) different numerical methods with step size, β„Ž = 10 seconds. Select the best numerical method to compute the temperature of the metal bar and justify your answer


1
Expert's answer
2021-12-08T04:32:01-0500

a)

"-\frac{dT}{T-T_m}=kdt"


"ln(T_m-T)=kt+c"


"T=T_m-e^{kt+c}=100-e^{kt+c}"


"T(0)=100-e^c=T_0"

"c=ln(100-T_0)"

"ln65\le c\le ln70"


"T(5)=100-e^{c+5k}=T_1"

"k=(ln(100-T_1)-c)\/5"

"(ln50-ln70)/5\le k\le (ln60-ln65)/5"

"-0.067\le k\le -0.016"


"100-70e^{-0.016t}\le T\le 100-65e^{-0.067t}"


reason for the selection of the method:

this is method to find exact value of variable


temperature of the metal bar after 100 seconds:

"100-70e^{-0.016\cdot100}\le T(100)\le 100-65e^{-0.067\cdot100}"

"85.87\le T(100)\le99.92"


b)

Euler methodΒ 

"T(0)=30"

"T'(t)=0.067(100-T)=f(t,T)"

"T_n=T_{n-1}+10f(t_{n-1},T_{n-1})"

"T_{10}=T(100)=99.8989\degree C"

error = "99.92-99.8989=0.0211"


Runge-Kutta 2 method:

"k_2=10f(t_0+10,T_0+k_1)"

"k_1=10f(t_0,T_0)=(10)f(0,30)=(10)\u22c5(4.69)=46.9"

"k_2=10f(t_0+10,T_0+k_1)=(10)f(10,76.9)=(10)\u22c5(1.5477)=15.477"

"T_1=T_0+(k_1+k_2)\/2=30+31.1885=61.1885"

"T(10)=61.1885"

........................

"T(100)=99.8078\degree C"

error = "99.92-99.8078=0.1122"


Runge-Kutta 3 method:

"k_1=10f(t_0,T_0)=(10)f(0,30)=(10)\u22c5(4.69)=46.9"

"k_2=10f(t_0+10\/2,T_0+k_1\/2)=(10)f(5,53.45)=(10)\u22c5(3.1188)=31.1885"

"k_3=10f(t_0+10,T_0+2k_2-k_1)=(10)f(10,45.477)=(10)\u22c5(3.653)=36.5304"

"T_1=T_0+(k_1+4k_2+k_3)\/6"

"T_1=30+[46.9+4(31.1885)+(36.5304)]\/6=64.6974"

"T(10)=64.6974"

...................................................

"T(100)=99.9255\degree C"

error = "99.9255-99.92=0.0055"


Minimal error is for Runge-Kutta 3 method, so this is the best numerical method to compute the temperature.